OpenStudy (anonymous):

Let y = f(z) = √4 − z2 and z = g(x) = 2x + 3. Compute the composition y = f(g(x)). Find the largest possible domain of x-values so that the composition y = f(g(x)) is defined. Need help.

OpenStudy (anonymous):

replace z in the function y with 2x + 3 to get the composition y = f(g(x))

OpenStudy (helder_edwin):

is it \[ \large f(z)=\sqrt{4}-z^2 \] or \[ \large f(z)=\sqrt{4-z^2} \]

OpenStudy (anonymous):

thanks